# PC-Gym: Benchmark Environments for Process Control Problems

### v2

@@ 2.4 Constrained Reinforcement Learning
In many real-world applications, particularly in chemical process control, the learning-based control problem must consider operational constraints. These constraints may represent safety limits, equipment capabilities, or product quality requirements. To incorporate these constraints, we can reformulate the learning-based control problem introduced in Section 2.2 as follows:
$\max_{\pi}\hskip 5.69054pt$ $\mathbb{E}_{\pi}\left[\sum^{T}_{t=0}\gamma^{t}r_{t}(X_{t},U_{t})\right]$
@@
where $g_{i}(\cdot)$ represents the $i$-th inequality constraint. The total number of constraints is denoted by $n_{g}$. Specifically, the problem imposes an individual chance constraint for each constraint function, given the uncertainty of the state. Explicitly, this allows for constraint violation with a given probability, $\alpha_{i,t}$. It is worth noting that one may also allow a probability for violation of the constraints jointly. This is known as a joint chance constraint. In practice, this is generally enabled by tuning the individual constraints imposed for example through constraint tightening.
This formulation leads us to the field of Constrained Reinforcement Learning (CRL). CRL aims to solve problems where the agent must maximize the expected return while satisfying constraints. Various approaches can be employed to solve constrained RL problems. These include Lagrangian methods, which transform the constrained optimization problem into an unconstrained one, reward shaping techniques that modify the reward function to incorporate penalties for constraint violations and safe policy optimization algorithms that directly optimize policies while ensuring constraint satisfaction throughout the learning process.
A key metric in evaluating constrained RL algorithms is the probability of constraint violation. Here, we consider the constraints jointly to provide a single metric to evaluate constraint satisfaction. Formally, for a given policy $\pi$ and constraint function $g_{i}(\cdot)$, the probability of constraint violation can be defined as:
@@
In practice, this probability is often estimated empirically according to an empirical cumulative distribution function defined as the result of multiple episodes or rollouts.
@@ 2.5 Policy Performance and Reproducibility Measures
Given the inherent stochastic nature of chemical systems, the same control policy $\pi$ can perform differently across runs, which we define as reproducibility [flageat2024expectedreturnaccountingpolicy]. Evaluating both the expected return $\mathbb{E}[J(\pi)|\pi]$ and reproducibility of policies is crucial. PC-Gym addresses this by reporting a performance measure, such as the median expected return $\tilde{J}_{\pi}$, along with a chosen dispersion metric across multiple training runs for chemical process control environments. This dispersion metric can be either the median absolute deviation $\text{MAD}(J(\pi))$ or the standard deviation $\sigma(J(\pi))$. For instance, the MAD quantifies policy reproducibility as:
$\text{MAD}(J(\pi))=\text{median}(|J(\pi)-\tilde{J}(\pi)|)$
Policies with high $\tilde{J}_{\pi}$ but poor reproducibility (high dispersion) may be unreliable in settings where consistent performance is critical, while those with slightly lower $\tilde{J}_{\pi}$ but good reproducibility (low dispersion) may be preferable for reliable operation. Reporting both the performance measure and a dispersion metric enables a more comprehensive and practically relevant evaluation of RL policies in chemical process control.
In addition to performance and reproducibility measures, the optimality gap provides a crucial metric for evaluating reinforcement learning policies in chemical process control. The optimality gap, $\Delta(\pi_{\theta})$, quantifies the difference between the performance of a learned policy $\pi_{\theta}$ and an optimal policy $\pi^{*}$:
$\Delta(\pi_{\theta})=J(\pi^{*})-J(\pi_{\theta})$ (11)
where $J(\pi_{\theta})$ and $J(\pi^{*})$ represent the expected returns of the learned and optimal policies, respectively. This metric offers a normalized measure of policy performance across different environments and indicates the potential for improvement. In chemical process control, where true optimal policies may be unknown, upper bounds or benchmark policies such as nonlinear model predictive control with a perfect model (NMPC) often serve as proxies for $\pi^{*}$.
Incorporating the optimality gap alongside performance measures like median expected return $\tilde{J}_{\pi}$ and dispersion metrics such as MAD or standard deviation provides a comprehensive evaluation framework. This approach balances absolute performance, consistency, and relative optimality, enabling a more nuanced assessment of RL policies in the stochastic and complex domain of chemical processes. Such an evaluation is essential for developing robust and reliable control strategies that meet the stringent requirements of industrial chemical applications.
@@ 2.6 Nonlinear MPC (Oracle)
To establish a performance benchmark for the reinforcement learning algorithms, we implemented a nonlinear Model Predictive Controller (NMPC) as an oracle. This controller leverages complete knowledge of the system dynamics, including perfect state estimation, to solve the optimal control problem at each time step. We define a normalization function for a vector z as $\bar{\textbf{z}}=\text{diag}(\textbf{z}_{max}-\textbf{z}_{min})^{-1}(\textbf{z}-\textbf{z}_{min})$. The NMPC is then formulated as a finite-horizon optimal control problem:
$\min_{\textbf{u}_{t},\ldots,\textbf{u}_{t+N-1}}J(\textbf{x},\textbf{u})$ $=\sum_{k=t}^{t+N}(\bar{\textbf{x}}_{k}-\bar{\textbf{x}}^{*})^{\top}\textbf{Q}(\bar{\textbf{x}}_{k}-\bar{\textbf{x}}^{*})+\sum_{k=t}^{t+N-1}(\bar{\textbf{u}}_{k}-\bar{\textbf{u}}_{k-1})^{\top}\textbf{R}(\bar{\textbf{u}}_{k}-\bar{\textbf{u}}_{k-1})$
@@
The MPC problem is set up with a user-defined horizon length $N$ and control weighting R. The objective function includes terms for both state tracking error and control effort. The state tracking error is computed as the squared difference between each controlled variable’s current state and setpoint. The control effort term is calculated using normalized input values to ensure consistent scaling across different input ranges.
@@ 3 Software Implementation
Central to PC-Gym is the make_env.py module, which serves as the environment constructor. Crucially, this module implements the Gymnasium [gym] interface, a widely adopted standard for reinforcement learning environments. By following the Gymnasium format, PC-Gym ensures compatibility with a broad range of existing RL algorithms and tools. The class accepts user-defined parameters as a dictionary, including model specifications, time steps, setpoints, and constraints.
@@ 4 Control Problems
This section describes four environments included in PC-Gym, indicating the types of dynamical systems included. The uncertainty in each control problem stems from measurement noise which given the perfect state estimator in the Oracle, results in no variation in the oracle’s state and control, and therefore cost trajectories. At the end of this section, all environments are benchmarked with three RL algorithms (SAC, PPO, and DDPG) at a setpoint point tracking control problem. All RL algorithms are trained using 50,000 timesteps using the default hyperparameters from Stable Baselines 3 [stable-baselines3] and the same seed. The choice of timestep budget was deemed large for this class of problems, allowing each algorithm to produce the highest performing policy for the same given seed. This policy is then fixed and evaluated on multiple instances of the environment. The following reward function follows the same notation as Oracle’s objective function as described in Section 2.6:
$r_{t}=-\left((\bar{\textbf{x}}_{t}-\bar{\textbf{x}}^{*})^{\top}\textbf{Q}(\bar{\textbf{x}}_{t}-\bar{\textbf{x}}^{*})+(\bar{\textbf{u}}_{k}-\bar{\textbf{u}}_{k-1})^{\top}\textbf{R}(\bar{\textbf{u}}_{k}-\bar{\textbf{u}}_{k-1})\right)$ (12)
where the state and control cost matrices, R and Q, are set to equal the corresponding matrix in the oracle for each case study.
@@ 4.3 Multistage Extraction Column
$\frac{dX_{n}}{dt}$ $=\frac{1}{V_{l}}(L(X_{n-1}-X_{n})-Q_{n})$ (15)
$\frac{dY_{n}}{dt}$ $=\frac{1}{V_{g}}(G(Y_{n+1}-Y_{n})+Q_{n})$ (16)
@@
$X_{n,eq}$ $=\frac{Y_{n}^{e}}{m}$ (17)
$Q_{n}$ $=K_{la}(X_{n}-X_{n,eq})V_{l}$ (18)
Here, $X_{n}$ and $Y_{n}$ are the solute concentrations in the liquid and gas phases, respectively, $L$ is the liquid flowrate, $G$ is the gas flowrate, $Q_{n}$ is the mass transfer rate, and $X_{n,eq}$ is the equilibrium concentration. The control variables are $L$ and $G$, while $X_{0}$ (feed concentration of liquid) and $Y_{6}$ (feed concentration of gas) are considered disturbances.
@@
$X_{0}$ Feed concentration of liquid (kg/m3) 0.60
$Y_{6}$ Feed concentration of gas (kg/m3) 0.050
@@ 4.3.1 Multistage Extraction Column Results
Figure 4: Comparison of reinforcement learning algorithms (SAC, PPO, DDPG) and an oracle for Multistage Extraction Column control. (a-b) Controlled variable response, (c-d) manipulated variable trajectory. Shaded areas show min-max ranges across simulations; solid lines represent median performance.
@@ 4.4.1 Crystallization Reactor Results
Figure 6: Comparison of reinforcement learning algorithms (SAC, PPO, DDPG) and an oracle for Crystalization Reactor control. (a) Controlled variable response, (b) manipulated variable trajectory, and (c) cumulative reward distribution. Shaded areas show ranges across simulations; solid lines represent median performance.
@@ 4.5.1 Four-Tank System Results
The performance of three reinforcement learning algorithms was evaluated on the Four Tank system control problem and compared to the Oracle controller. The oracle MPC used a prediction horizon of 17 steps, an identity Q matrix, and a zero R matrix. An episode was defined as 1000 seconds with 60 timesteps. Figure 7 illustrates the control performance of these algorithms, showing both controlled variables ($h1$, $h2$) and manipulated variables ($V_{1}$, $V_{2}$), while Figure 8 displays the distribution of cumulative rewards.
@@ 4.6 Overview of Case Studies
The optimality gap, as previously discussed in Section 2.5, holds significant importance in reinforcement learning for several reasons. Primarily, it offers a scalar measure that quantifies how far a learned policy is from optimality, enabling direct comparisons between different algorithms. For our study in PC-Gym environments, we calculate the empirical optimality gap with normalised cumulative rewards, and with the optimal policy approximated by the oracle $\pi^{*}\approx\pi_{o}$:
$\hat{\Delta}(\pi_{\theta})=\sum_{i=1}^{M}J^{(i)}(\pi_{o})-\sum_{i=1}^{M}J^{(i)}(\pi_{\theta})$ (32)
where $M$ is the number of episodes, and $r^{(i)}_{t}(\pi)$ is the reward at time $t$ in episode $i$ under policy $\pi$ where $\pi_{o}$ represents the oracle and $\pi_{\theta}$ represents an RL policy either optimized with SAC, PPO, or DDPG.
@@ 6.3 Experimental Results
The constrained agent successfully learned to maintain the temperature within the specified bounds, as evident from the temperature trajectory in Figure 11. In contrast, the base agent violated these limits in pursuit of optimal setpoint tracking for the concentration (Figure 11), especially in the overshoot of both setpoints. To quantify constraint satisfaction, we leverage the empirical cumulative probability of constraint violation, $V$:
@@ 7 Reward Functions
$r_{t}=-\lvert|\bar{\textbf{x}}^{*}-\bar{\textbf{x}}\rvert|_{1}$ (37)
@@
$r_{t}=-\lvert|\bar{\textbf{x}}^{*}-\bar{\textbf{x}}\rvert|_{2}$ (38)
@@
$r(t)=\begin{cases}1,&\text{if }-\lvert|\bar{\textbf{x}}^{*}-\bar{\textbf{x}}\rvert|_{2}<\epsilon\\ 0,&\text{otherwise}\end{cases}$ (39)
@@ 9 Acknowledgements
Maximilian Bloor would like to acknowledge funding provided by the Engineering & Physical Sciences Research Council, United Kingdom, through grant code EP/W524323/1. Calvin Tsay acknowledges support from a BASF/Royal Academy of Engineering Senior Research Fellowship. José Torraca would like to acknowledge funding provided by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001.

### v4

@@ 2.4 Constrained Reinforcement Learning
In many real-world applications, particularly in chemical process control, the learning-based control problem must consider operational constraints. These constraints may represent safety limits, equipment capabilities, or product quality requirements. To incorporate these constraints, we can reformulate the learning-based control problem introduced in Section 2.2 into an example constrained reinforcement learning formulation:
$\max_{\pi}\hskip 5.69054pt$ $\mathbb{E}_{\pi}\left[\sum^{T-1}_{t=0}\gamma^{t}r_{t+1}(X_{t},U_{t})\right]$
@@
where $g_{i}(\cdot)$ represents the $i$-th inequality constraint. The total number of constraints imposed over the time horizon is $n_{g}\times T$, therefore the problem imposes an individual chance constraint for each constraint function at each time index, given the uncertainty of the state over the time horizon. Explicitly, this allows for constraint violation with a given probability, $\alpha_{i,t}$. It is worth noting that one may also allow a probability for violation of the constraints jointly. This is known as a joint chance constraint [paulson_smpc, PETSAGKOURAKIS202235]. In practice, this is generally enabled by tuning the individual constraints imposed, for example through constraint tightening [PAN2021107462]. Recent work has shown that constraint-tightening approaches can effectively guarantee satisfaction of joint chance constraints with high probability while maintaining good control performance.
Various approaches can be employed to solve constrained RL problems. These include: Lagrangian methods [BORKAR2005207, peng_separated_2021], which transform the constrained optimization problem into an unconstrained one by incorporating constraints into the objective function through Lagrange multipliers; reward shaping techniques that modify the reward function to incorporate penalties for constraint violations [peng_separated_2021]; and safe policy optimization algorithms like Constrained Policy Optimization (CPO) that directly optimize policies while ensuring constraint satisfaction throughout the learning process [pmlr-v70-achiam17a]. The Separated Proportional-Integral Lagrangian (SPIL) approach combines the benefits of both penalty and Lagrangian methods to achieve better performance while maintaining safety guarantees [peng_separated_2021].
A key metric in evaluating constrained RL algorithms is the joint probability of constraint violation. Here, the constraints are treated jointly to provide a single metric to evaluate constraint satisfaction. Formally, for a given policy $\pi$ the joint probability of constraint violation can be defined as:
@@
In practice, this probability is often estimated empirically according to an empirical cumulative distribution function defined as the result of multiple episodes or rollouts [petsagkourakis2022chance].
@@ 2.5 Policy Performance and Reproducibility Measures
Given the inherent stochastic nature of chemical systems and RL policies, the same control policy $\pi$ can perform differently across runs, which we define as reproducibility [flageat2024expectedreturnaccountingpolicy]. Evaluating both the expected return $J(\pi)$ and reproducibility of policies is crucial. PC-Gym addresses this by reporting a performance measure, such as the median return $\tilde{J}(\pi)$, along with a chosen dispersion metric across multiple repetitions in a fixed chemical process control environment. This dispersion metric can either be the median absolute deviation $\text{MAD}(\pi)$ or the standard deviation $J_{\sigma}(\pi)$ of returns. For instance, $\text{MAD}(\pi)$ quantifies policy reproducibility as:
$\text{MAD}(\pi)=\text{Median}_{\pi}\left[\left\lvert\sum_{t=0}^{T-1}\gamma^{t}r_{t+1}(X_{t},U_{t})-\tilde{J}(\pi)\right\rvert\right]\,.$ (11)
In practice, this quantity is estimated from rolling out the policy over a number of episodesin the fixed environment. Policies with high $\tilde{J}(\pi)$ but high dispersion may be unreliable in settings where consistent performance is critical, while those with slightly lower $\tilde{J}(\pi)$ but low dispersion may be preferable for reliable operation. Reporting both the performance measure and a dispersion metric enables a more comprehensive and practically relevant evaluation of RL policies in chemical process control.
In addition to performance and reproducibility measures, the optimality gap provides a crucial metric for evaluating reinforcement learning policies in chemical process control. The optimality gap, $\Delta(\pi_{\phi})$, quantifies the difference between the performance of a learned policy $\pi_{\phi}$ and an optimal policy $\pi^{*}$:
$\Delta(\pi_{\phi})=J(\pi^{*})-J(\pi_{\phi})$ (12)
where $J(\pi_{\phi})$ and $J(\pi^{*})$ represent the expected returns of the learned and optimal policies, respectively. This metric offers a normalized measure of policy performance across different environments and indicates the potential for improvement. In chemical process control, where true optimal policies may be unknown, upper bounds or benchmark policies such as nonlinear model predictive control with a perfect model (NMPC) often serve as proxies for $\pi^{*}$.
Incorporating the optimality gap alongside performance measures like median return $\tilde{J}({\pi})$ and dispersion metrics such as MAD or standard deviation provides a comprehensive evaluation framework. This approach balances absolute performance, consistency, and relative optimality, enabling a more nuanced assessment of RL policies in the stochastic and complex domain of chemical processes. Such an evaluation is essential for developing robust and reliable control strategies that meet the stringent requirements of industrial chemical applications.
@@ 2.6 Nonlinear MPC (Oracle)
To establish a performance benchmark for the RL algorithms, we implemented a nonlinear Model Predictive Controller (NMPC) as an oracle. This controller leverages complete knowledge of the system dynamics, including perfect state estimation, to solve the optimal control problem at each time step. We define a normalization function for a vector z as $\bar{\textbf{z}}=\text{diag}(\textbf{z}_{max}-\textbf{z}_{min})^{-1}(\textbf{z}-\textbf{z}_{min})$. The NMPC is then formulated as a finite-horizon optimal control problem:
$\min_{\textbf{u}_{t},\ldots,\textbf{u}_{t+N-1}}J(\textbf{x}_{t},\textbf{u}_{t})$ $=\sum_{k=t+1}^{t+N}(\bar{\textbf{x}}_{k}-\bar{\textbf{x}}^{*})^{\top}\textbf{Q}(\bar{\textbf{x}}_{k}-\bar{\textbf{x}}^{*})+\sum_{k=t}^{t+N-1}(\bar{\textbf{u}}_{k}-\bar{\textbf{u}}_{k-1})^{\top}\textbf{R}(\bar{\textbf{u}}_{k}-\bar{\textbf{u}}_{k-1})$
@@
The NMPC problem is set up with a user-defined horizon length $N$ and control weighting R. The objective function includes terms for both state tracking error and control effort. The state tracking error is computed as the squared difference between each controlled variable’s current state and setpoint. The control effort term is calculated using normalized input values to ensure consistent scaling across different input ranges.
@@ 3 Software Implementation
Central to PC-Gym is the make_env.py module, which serves as the environment constructor. This module implements the Gymnasium [gym] interface, a widely adopted standard for reinforcement learning environments. By following the Gymnasium format, PC-Gym ensures compatibility with a broad range of existing RL algorithms and tools. The class accepts user-defined parameters as a dictionary, including model specifications, time steps, setpoints, and constraints.
@@ 4 Control Problems
This section describes four environments included in PC-Gym, indicating the types of dynamical systems included. The uncertainty in each control problem stems from measurement noise which given the perfect state estimator in the Oracle, results in no variation in the oracle’s state and control, and therefore cost trajectories. At the end of this section, all environments are benchmarked with three RL algorithms (SAC, PPO, and DDPG) at a setpoint point tracking control problem. All RL algorithms are trained using 50,000 timesteps using the default hyperparameters from Stable Baselines 3 [stable-baselines3] and the same seed. The choice of timestep budget was deemed large for this class of problems, allowing each algorithm to produce the highest performing policy for the same given seed. This policy is then fixed and evaluated on multiple instances of the environment. The following reward function follows the same notation as the Oracle’s objective function as described in Section 2.6:
$r_{t+1}(\textbf{x}_{t},\textbf{u}_{t})=-\left((\bar{\textbf{x}}_{t+1}-\bar{\textbf{x}}^{*})^{\top}\textbf{Q}(\bar{\textbf{x}}_{t+1}-\bar{\textbf{x}}^{*})+(\bar{\textbf{u}}_{t}-\bar{\textbf{u}}_{t-1})^{\top}\textbf{R}(\bar{\textbf{u}}_{t}-\bar{\textbf{u}}_{t-1})\right)$ (13)
where the state and control cost matrices, R and Q, are set to equal the corresponding matrix in the oracle for each case study, and $\textbf{x}_{t+1}$ is obtained through forward simulation of the state space model given the current state and control pair, $(\textbf{x}_{t},\textbf{u}_{t})$ and realization of uncertain parameters.
@@ 4.3 Multistage Extraction Column
$\frac{dC_{X_{n}}}{dt}$ $=\frac{1}{V_{l}}(L(C_{X_{n-1}}-C_{X_{n}})-F_{n})$ (16)
$\frac{dC_{Y_{n}}}{dt}$ $=\frac{1}{V_{g}}(G(C_{Y_{n+1}}-C_{Y_{n}})+F_{n})$ (17)
@@
$C_{X_{n,eq}}$ $=\frac{C_{Y_{n,eq}}}{m}$ (18)
$F_{n}$ $=K_{la}(C_{X_{n}}-C_{X_{n,eq}})V_{l}$ (19)
Here, $C_{X_{n}}$ and $C_{Y_{n}}$ are the solute concentrations in the liquid and gas phases, respectively, $L$ is the liquid flowrate, $G$ is the gas flowrate, $F_{n}$ is the mass transfer rate, and $C_{X_{n,eq}}$ is the equilibrium concentration. The control variables are $L$ and $G$, while $C_{X_{0}}$ (feed concentration of liquid) and $C_{Y_{6}}$ (feed concentration of gas) are considered disturbances.
@@
$C_{X_{0}}$ Feed concentration of liquid (kg/m3) 0.60
$C_{Y_{6}}$ Feed concentration of gas (kg/m3) 0.050
@@ 4.3.1 Multistage Extraction Column Results
Figure 4: Comparison of reinforcement learning algorithms (SAC, PPO, DDPG) and an oracle for Multistage Extraction Column control. Controlled variable response, manipulated variable trajectory. Shaded areas show min-max ranges across simulations; solid lines represent median performance.
@@ 4.4.1 Crystallization Reactor Results
Figure 6: Comparison of reinforcement learning algorithms (SAC, PPO, DDPG) and an oracle for Crystalization Reactor control. Controlled variable response, manipulated variable trajectory, and cumulative reward distribution. Shaded areas show ranges across simulations; solid lines represent median performance.
@@ 4.5.1 Four-Tank System Results
The performance of three reinforcement learning algorithms was evaluated on the Four Tank system control problem and compared to the Oracle controller. The oracle MPC used a prediction horizon of 17 steps, an identity Q matrix, and a zero R matrix. An episode was defined as 1000 seconds with 60 timesteps. Figure 7 illustrates the control performance of these algorithms, showing both controlled variables ($h_{1}$, $h_{2}$) and manipulated variables ($V_{1}$, $V_{2}$), while Figure 8 displays the distribution of cumulative rewards.
@@ 4.6 Overview of Case Studies
The optimality gap, as previously discussed in Section 2.5, holds significant importance in reinforcement learning for several reasons. Primarily, it offers a scalar measure that quantifies how far a learned policy is from optimality, enabling direct comparisons between different algorithms. For our study in PC-Gym environments, we calculate the empirical optimality gap with normalized cumulative rewards, and with the optimal policy approximated by the oracle $\pi^{*}\approx\pi_{o}$:
$\hat{\Delta}(\pi_{\phi})=\tilde{J}(\pi_{o})-\tilde{J}(\pi_{\phi})$ (33)
where $M$ is the number of episodes, and $\tilde{J}(\pi)$ is the median return under policy $\pi$ where $\pi_{o}$ represents the oracle and $\pi_{\phi}$ represents an RL policy either optimized with SAC, PPO, or DDPG. It should be noted that $\tilde{J}(\cdot)$ uses normalized return values to allow for comparison between case studies.
@@ 6.3 Experimental Results
The constrained agent successfully learned to maintain the temperature within the specified bounds, as evident from the temperature trajectory in Figure 11. In contrast, the base agent violated these limits in pursuit of optimal setpoint tracking for the concentration (Figure 11), especially in the overshoot of both setpoints. To quantify constraint satisfaction, we leverage the empirical cumulative probability of constraint violation:
@@ 7 Reward Functions
$r_{t+1}(\textbf{x}_{t},\textbf{u}_{t})=-\lvert|\bar{\textbf{x}}^{*}-\bar{\textbf{x}}_{t+1}\rvert|_{1}$ (38)
@@
$r_{t+1}(\textbf{x}_{t},\textbf{u}_{t})=-\lvert|\bar{\textbf{x}}^{*}-\bar{\textbf{x}}_{t+1}\rvert|_{2}$ (39)
@@
$r_{t+1}(\textbf{x}_{t},\textbf{u}_{t})=\begin{cases}1,&\text{if }-\lvert|\bar{\textbf{x}}^{*}-\bar{\textbf{x}}_{t+1}\rvert|_{2}<\epsilon\\ 0,&\text{otherwise}\,.\end{cases}$ (40)
In all cases, the next state $\bar{\textbf{x}}_{t+1}$ is obtained through forward simulation of the state space model, given a realization of uncertain parameters and the current state-control pair.
@@ 9 Acknowledgements
Maximilian Bloor would like to acknowledge funding provided by the Engineering & Physical Sciences Research Council, United Kingdom, through grant code EP/W524323/1. Calvin Tsay acknowledges support from a BASF/Royal Academy of Engineering Senior Research Fellowship. José Torraca would like to acknowledge funding provided by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. The authors would like to acknowledge Panagiotis Petsagkourakis for productive discussions over the course of this work.